# PUNCH: Physics-informed Uncertainty-aware Network for Coronary Hemodynamics

### v1

@@ Abstract
Coronary microvascular dysfunction (CMD) affects millions worldwide yet remains underdiagnosed because gold-standard physiological measurements are invasive and variably reproducible. We introduce a non-invasive, uncertainty-aware framework for estimating coronary flow reserve (CFR) directly from standard angiography. The system integrates physics-informed neural networks with variational inference to infer coronary blood flow from first-principles models of contrast transport, without requiring ground-truth flow measurements. The pipeline runs in approximately three minutes per patient on a single GPU, with no population-level training.
Using 1,000 synthetic spatiotemporal intensity maps (kymographs) with controlled noise and artifacts, the framework reliably identifies degraded data and outputs appropriately inflated uncertainty estimates, showing strong correspondence between predictive uncertainty and error (Pearson $r=0.997$, Spearman $\rho=0.998$). Clinical validation in 12 patients shows strong agreement between PUNCH-derived CFR and invasive bolus thermodilution (Pearson $r=0.90$, $p=6.3\times 10^{-5}$). We focus on the LAD, the artery most commonly assessed in routine CMD testing. Probabilistic CFR estimates have confidence intervals narrower than the variability of repeated invasive measurements.
By transforming routine angiography into quantitative, uncertainty-aware assessment, this approach enables scalable, safer, and more reproducible evaluation of coronary microvascular function. Because standard angiography is widely available globally, the framework could expand access to CMD diagnosis and establish a new paradigm for physics-informed, patient-specific inference from clinical imaging.
@@ Introduction
Coronary microvascular dysfunction (CMD) is a prevalent yet underdiagnosed condition, with evidence of microvascular dysfunction reported in up to 70% of patients with angina and non-obstructive coronary arteries undergoing invasive angiography Kunadian et al. (2020). This diagnostic oversight disproportionately impacts female populations Ong and others (2018). Unlike epicardial stenosis, CMD reflects abnormalities in microvascular flow and structure that are invisible on standard angiography but critically influence myocardial perfusion and patient outcomes Murthy and others (2014); Taqueti and Di Carli (2018).
The gold standard for diagnosing coronary microvascular function is the measurement of coronary flow reserve (CFR) using invasive catheter-based techniques such as Doppler flow wires or thermodilution Taqueti and Di Carli (2018); Barbato et al. (2014). In thermodilution, a known volume of cold saline is injected into the coronary artery, and downstream temperature changes are recorded by a thermistor-tipped catheter. The resulting temperature–time curve—an indicator-dilution profile—allows for estimation of blood flow based on the area under the curve Fearon and others (2003). Although reliable, thermodilution requires precise timing, is sensitive to injection speed and temperature variability, and carries procedural risks inherent to placement of a coronary guide-wire in the distal coronary artery De Bruyne et al. (2001); Fearon and others (2003). These practical constraints have limited the widespread evaluation of CFR, leaving most patients without a definitive functional diagnosis.
Figure 1: PUNCH enables uncertainty-aware estimation of coronary flow reserve from standard angiography. Fluoroscopy images acquired at rest and hyperemia undergo vessel centerline tracking to generate spatiotemporal intensity maps (kymographs). These kymographs serve as input to a physics-informed neural network that jointly optimizes data fidelity ($\mathcal{L}_{Data}$), advection-diffusion physics constraints ($\mathcal{L}_{PDE}$), and variational inference regularization ($\mathcal{L}_{KL}$). The framework outputs posterior distributions of rest velocity, hyperemic velocity, and coronary flow reserve (CFR), providing probabilistic estimates with quantified uncertainty. The dual-branch architecture models both physiological states while sharing a common latent variable z to capture measurement uncertainty from noise, motion artifacts, and incomplete contrast filling.
@@
Here, we introduce an uncertainty-aware, physics-informed framework to estimate coronary flow reserve (CFR) directly from standard coronary angiography. The key insight is simple: contrast material injected during angiography behaves as an intravascular tracer. As it travels through the coronary artery, it is transported by blood flow and simultaneously spreads due to velocity heterogeneity and mixing. These observable spatiotemporal patterns encode information about coronary blood flow and microvascular resistance.
@@
The goal of the proposed framework is to estimate the coronary flow reserve (CFR) from routine coronary angiography while providing an explicit measure of confidence in each prediction. CFR is defined clinically as the ratio of blood flow during pharmacologically induced hyperemia to that at rest and is a key marker of coronary microvascular function.
@@
The correlation between predicted and invasively measured coronary flow reserve (CFR) was assessed using Pearson’s correlation coefficient ($r$). Agreement between methods was evaluated using Bland–Altman analysis Bland and Altman (1986) and the concordance correlation coefficient (CCC). Calibration of predicted uncertainty intervals was assessed by computing the fraction of ground-truth (synthetic) or invasive (clinical) CFR values contained within the model’s 95% confidence intervals. For synthetic datasets where true values were known, the uncertainty–error correlation was calculated to quantify whether higher predicted uncertainty corresponded to greater absolute error.
@@
The propagation of contrast along the vessel centerline was modeled as a one-dimensional transport process governed by advection and dispersion, reflecting contrast motion with blood flow and its spreading due to velocity heterogeneity and mixing. Specifically, contrast intensity $I(s,t)$ was generated by solving eq. (3). Physiologically plausible velocity profiles were randomly generated for each sequence. Mean resting velocities were sampled between $10$ and $20\frac{cm}{s}$ for typical cases, while 25% of the sequences were designated as atypical stress-test cases with elevated resting velocities of $20–40\frac{cm}{s}$. Spatial velocity variations were introduced using smooth, localized modulations along the vessel to mimic focal acceleration or deceleration. Hyperemic flow was simulated by scaling resting velocities using a randomly sampled coronary flow reserve (CFR), drawn from 1.5 to 4.0 for typical cases and 1.0 to 2.0 for low-CFR cases.
@@ Results
Clinical validation was performed in a cohort of 12 patients who underwent a standard invasive coronary flow reserve (CFR) assessment using the bolus thermodilution technique. For each patient, three repeated measurements were acquired at rest and during pharmacologically induced hyperemia. Invasive CFR was calculated as the ratio of mean hyperemic to mean resting flow values Barbato et al. (2014); Fearon and others (2003). The CFR estimates derived from PUNCH were obtained independently for each case and summarized by their posterior mean and standard deviation.
@@
The PUNCH estimates demonstrated a strong association with invasive measurements, with a Pearson correlation of $r=0.90$ ($p=6.3\times 10^{-5}$) and a Spearman rank correlation of $\rho=0.76$ ($p=0.004$). Ordinary least squares (OLS) regression yielded an $R^{2}$ of 0.81, indicating that a substantial proportion of the variance in invasive CFR was explained by the PUNCH predictions (Fig. 5B).
@@
Figure 5: PUNCH demonstrates strong agreement with invasive CFR assessment. (a) Input kymograph (blue) to PUNCH, output kymograph (yellow) and overlay of input and output kymograph (b) Correlation plot comparing PUNCH-derived CFR posterior means against invasive bolus thermodilution measurements in 12 patients shows strong linear association (Pearson $r=0.90$, $p=6.3\times 10^{-5}$; Deming regression slope =1.01=1.01, intercept =0.49=-0.49). Error bars represent PUNCH posterior standard deviations and invasive measurement variability across three repeated acquisitions. (c) Bland–Altman analysis reveals mean bias of 0.45-0.45 (SD $\pm 0.75$) with 95% limits of agreement [1.91,1.02][-1.91,1.02], indicating modest systematic underestimation. All 12 cases fall within agreement bounds. (d) Individual patient predictions with 95% credible intervals (shaded regions) demonstrate consistency between predicted uncertainty and observed invasive values, with narrower PUNCH confidence intervals than intra-patient invasive variability, suggesting improved reproducibility.
@@ Discussion
We present a physics-informed, uncertainty-aware framework for non-invasive diagnosis of coronary microvascular dysfunction (CMD) that transforms routine angiography into a functional assessment tool. Our approach integrates (1) physics-based learning directly from unlabeled angiographic sequences with per-patient inference; (2) variational inference that produces calibrated uncertainty estimates tied to image quality; and (3) a dual-branch architecture that jointly models resting and hyperemic states to produce physiologically consistent coronary flow reserve (CFR) confidence intervals. By extracting quantitative physiology from standard-of-care imaging, this framework addresses a critical unmet need in cardiovascular diagnostics. Importantly, the complete analysis pipeline executes in approximately three minutes per patient on a single NVIDIA A40 GPU, with no requirement for prior population-level training, enabling rapid, on-demand physiological assessment directly at the point of care.
@@
Comparison to Existing Paradigms. Computational fluid dynamics (CFD)-based methods require high-resolution CT imaging, three-dimensional vessel reconstruction, and extensive computation (often hours per patient), fundamentally limiting clinical scalability Taylor et al. (2013); Morris et al. (2016). Supervised machine learning approaches require large labeled datasets that simply do not exist for challenging phenotypes such as CMD, where invasive ground truth measurements are available only in small registries Itu et al. (2016); Kweon et al. (2020). In contrast, our framework learns directly from routinely acquired angiograms without labeled flow data, computes patient-specific solutions in minutes, and provides probabilistic estimates with quantified uncertainty; capabilities absent in prior approaches. The per-patient training paradigm eliminates concerns about generalization between heterogeneous populations or data leakage between training and test sets. In particular, our predicted CFR confidence intervals were narrower than the intra-patient variability observed in repeated invasive thermodilution measurements (median standard deviation approximately 12%) Barbato et al. (2014), indicating both precision and potential for improved reproducibility relative to the current invasive gold standard.
@@
Sukirt Thakur, Yang Zhou, Srinivas Paruchuri, and Scott Burger are employees of and hold equity interests in AngioInsight, Inc., a company developing artificial intelligence solutions in cardiac health. Brahmajee Nallamothu and Alberto Figueroa are employed by the University of Michigan and hold an equity interest in AngioInsight, Inc. as consultants and co-founders.

### v2

@@ Abstract
More than 10 million coronary angiograms are performed globally each year, providing a gold standard for detecting obstructive coronary artery disease. Yet, no obstructive lesions are identified in 70% of patients evaluated for ischemic heart disease. Up to half of these patients have undiagnosed, life-limiting coronary microvascular dysfunction (CMD), which remains under-detected due to the limited availability of invasive tools required to measure coronary flow reserve (CFR). Here, we introduce PUNCH, a non-invasive, uncertainty-aware framework for estimating CFR directly from standard coronary angiography. PUNCH integrates physics-informed neural networks with variational inference to infer coronary blood flow from first-principles models of contrast transport, without requiring ground-truth flow measurements or population-level training. The pipeline runs in approximately three minutes per patient on a single GPU. Validated on synthetic angiograms with controlled noise and imaging artifacts, as well as on clinical bolus thermodilution data from 20 patients, PUNCH demonstrates accurate and uncertainty-calibrated CFR estimation. This approach establishes a new paradigm for CMD diagnosis and illustrates how physics-informed inference can substantially expand the diagnostic utility of available clinical imaging.
@@ Introduction
Coronary microvascular dysfunction (CMD) is a prevalent yet underdiagnosed condition, with evidence of microvascular dysfunction reported in up to 70% of patients with angina and non-obstructive coronary arteries undergoing invasive angiography Kunadian et al. (2020). This diagnostic oversight disproportionately impacts female populations Ong and others (2018). Unlike epicardial stenosis, CMD reflects abnormalities in flow and structure that are invisible on standard angiography but critically influence myocardial perfusion and patient outcomes Murthy and others (2014); Taqueti and Di Carli (2018).
The gold standard for diagnosing coronary microvascular function is the measurement of coronary flow reserve (CFR) using invasive catheter-based techniques such as Doppler flow wires or thermodilution Taqueti and Di Carli (2018); Barbato et al. (2014). In thermodilution, a known volume of cold saline is injected into the coronary artery, and downstream temperature changes are recorded by a thermistor-tipped catheter. The resulting temperature–time curve—an indicator-dilution profile—allows for estimation of blood flow based on the area under the curve Fearon and others (2003). Although reliable, thermodilution requires precise timing, is sensitive to injection technique and temperature variability, and carries procedural risks inherent to placement of a coronary guide-wire in the distal coronary artery De Bruyne et al. (2001); Fearon and others (2003). These practical constraints have limited the widespread invasive evaluation of CFR, leaving most patients without a definitive functional diagnosis.
Figure 1: PUNCH enables uncertainty-aware estimation of CFR from standard angiography. Fluoroscopy images acquired at rest and hyperemia undergo vessel centerline tracking to generate spatiotemporal intensity maps (kymographs). These kymographs serve as input to a physics-informed neural network that jointly optimizes data fidelity ($\mathcal{L}_{Data}$), advection-diffusion physics constraints ($\mathcal{L}_{PDE}$), and variational inference regularization ($\mathcal{L}_{KL}$). The framework outputs posterior distributions of rest velocity, hyperemic velocity, and CFR, providing probabilistic estimates with quantified uncertainty. The dual-branch architecture models both physiological states while sharing a common latent variable z to capture measurement uncertainty from noise, motion artifacts, and incomplete contrast filling.
@@
Here, we introduce an uncertainty-aware, physics-informed framework to estimate CFR directly from standard coronary angiography. The key insight is simple: contrast material injected during angiography behaves as an intravascular tracer. As it travels through the coronary artery, it is transported by blood flow and simultaneously spreads due to velocity heterogeneity and mixing. These observable spatiotemporal patterns encode information about coronary blood flow and microvascular resistance.
@@
The goal of the proposed framework is to estimate the CFR from routine coronary angiography while providing an explicit measure of confidence in each prediction. CFR is defined clinically as the ratio of blood flow during pharmacologically induced hyperemia to that at rest and is a key marker of coronary microvascular function.
@@
The correlation between predicted and invasively measured CFR was assessed using Spearman rank correlation. Agreement between methods was evaluated using Bland–Altman analysis Bland and Altman (1986) and the concordance correlation coefficient (CCC). Calibration of predicted uncertainty intervals was assessed by computing the fraction of ground-truth (synthetic) or invasive (clinical) CFR values contained within the model’s 95% confidence intervals. For synthetic datasets where true values were known, the uncertainty–error correlation was calculated to quantify whether higher predicted uncertainty corresponded to greater absolute error.
@@
The propagation of contrast along the vessel centerline was modeled as a one-dimensional transport process governed by advection and dispersion, reflecting contrast motion with blood flow and its spreading due to velocity heterogeneity and mixing. Specifically, contrast intensity $I(s,t)$ was generated by solving eq. (3). Physiologically plausible velocity profiles were randomly generated for each sequence. Mean resting velocities were sampled between $10$ and $20\frac{cm}{s}$ for typical cases, while 25% of the sequences were designated as atypical stress-test cases with elevated resting velocities of $20–40\frac{cm}{s}$. Spatial velocity variations were introduced using smooth, localized modulations along the vessel to mimic focal acceleration or deceleration. Hyperemic flow was simulated by scaling resting velocities using a randomly sampled CFR, drawn from 1.5 to 4.0 for typical cases and 1.0 to 2.0 for low-CFR cases.
@@ Results
Clinical validation was performed in a cohort of 12 patients who underwent a standard invasive CFR assessment using the bolus thermodilution technique. For each patient, three repeated measurements were acquired at rest and during pharmacologically induced hyperemia using intravenous adenosine at 140 mcg/kg/min. Invasive CFR was calculated as the ratio of mean hyperemic to mean resting flow values Barbato et al. (2014); Fearon and others (2003). The CFR estimates derived from PUNCH were obtained independently for each case and summarized by their posterior mean and standard deviation.
@@
The PUNCH estimates demonstrated a strong association with invasive measurements, with a Spearman rank correlation of $\rho=0.76$ ($p=0.004$). Ordinary least squares (OLS) regression yielded an $R^{2}$ of 0.81, indicating that a substantial proportion of the variance in invasive CFR was explained by the PUNCH predictions (Fig. 5B).
@@
Figure 5: PUNCH demonstrates strong agreement with invasive CFR assessment. (a) Input kymograph (blue) to PUNCH, output kymograph (yellow) and overlay of input and output kymograph (b) Correlation plot comparing PUNCH-derived CFR posterior means against invasive bolus thermodilution measurements in 12 patients shows strong linear association (a Spearman rank correlation of $\rho=0.76$ ($p=0.004$); Deming regression slope =1.01=1.01, intercept =0.49=-0.49). (c) Bland–Altman analysis reveals mean bias of 0.45-0.45 (SD $\pm 0.75$) with 95% limits of agreement [1.91,1.02][-1.91,1.02], indicating modest systematic underestimation. All 12 cases fall within agreement bounds. (d) Individual patient predictions with 95% credible intervals (shaded regions) demonstrate consistency between predicted uncertainty and observed invasive values, with narrower PUNCH confidence intervals than intra-patient invasive variability, suggesting improved reproducibility. Error bars represent PUNCH posterior standard deviations and invasive measurement variability across three repeated acquisitions.
@@ Discussion
We present a physics-informed, uncertainty-aware framework for non-invasive diagnosis of coronary microvascular dysfunction (CMD) that transforms routine angiography into a functional assessment tool. Our approach integrates (1) physics-based learning directly from unlabeled angiographic sequences with per-patient inference; (2) variational inference that produces calibrated uncertainty estimates tied to image quality; and (3) a dual-branch architecture that jointly models resting and hyperemic states to produce physiologically consistent CFR confidence intervals. By extracting quantitative physiology from standard-of-care imaging, this framework addresses a critical unmet need in cardiovascular diagnostics. Importantly, the complete analysis pipeline executes in approximately three minutes per patient on a single NVIDIA A40 GPU, with no requirement for prior population-level training, enabling rapid, on-demand physiological assessment directly at the point of care.
@@
Comparison to Existing Paradigms. Computational fluid dynamics (CFD)-based methods require high-resolution CT imaging, three-dimensional vessel reconstruction, and extensive computation (often hours per patient), fundamentally limiting clinical scalability Taylor et al. (2013); Morris et al. (2016). Supervised machine learning approaches require large labeled datasets that simply do not exist for challenging phenotypes such as CMD, where invasive ground truth measurements are available only in small registries Itu et al. (2016); Kweon et al. (2020). In contrast, our framework learns directly from routinely acquired angiograms without labeled flow data, computes patient-specific solutions in minutes, and provides probabilistic estimates with quantified uncertainty; capabilities absent in prior approaches. The per-patient training paradigm eliminates concerns about generalization between heterogeneous populations or data leakage between training and test sets. In particular, our predicted CFR confidence intervals were narrower than the intra-patient variability observed in repeated invasive thermodilution measurements (median standard deviation approximately 12%) Barbato et al. (2014), indicating both precision and potential for improved reproducibility relative to the current invasive methods.
@@
Sukirt Thakur, Yang Zhou, Srinivas Paruchuri, and Scott Burger are employees of and hold equity interests in AngioInsight, Inc., a company developing artificial intelligence solutions in cardiac health. Brahmajee K. Nallamothu and C. Alberto Figueroa are employed by the University of Michigan and hold an equity interest in AngioInsight, Inc. as consultants and co-founders.